Suppose that you determine the density of a mineral by measuring its mass (m) (4.635±0.002 g) and its volume (1.13±0.05 mL). What is the uncertainty in the computed density?
Answer #1

Density(d) = mass (m) /volume(v)

d = 4.635÷1.13 = 4.101 g/ml.

Uncertainty in density

d = d× [( m/m)2 + (v/v)2] 1/2

= 4.101× [ 0.002÷4.635)2 + (0.05/1.13)2]1/2

= 4.101×[ 0.000000186 + 0.00195786671]1/2

= 4.101×[0.001958]1/2

= 4.101×0.0444

= 0.182

So, uncertainty in the computed density = 0.182